# NormAd: A Framework for Measuring the Cultural Adaptability of Large Language Models

### v6

@@ Abstract
To be effectively and safely deployed to global user populations, large language models (LLMs) must adapt outputs to user values and culture, not just know about them. We introduce NormAd, an evaluation framework to assess LLMs’ cultural adaptability, specifically measuring their ability to judge social acceptability across different levels of cultural norm specificity, from abstract values to explicit social norms. As an instantiation of our framework, we create NormAd-Eti, a benchmark of 2.6k situational descriptions representing social-etiquette related cultural norms from 75 countries. Through comprehensive experiments on NormAd-Eti, we find that LLMs struggle to accurately judge social acceptability across these varying degrees of cultural contexts and show stronger adaptability to English-centric cultures over those from the Global South. Even in the simplest setting where the relevant social norms are provided, our best models’ performance ($<$82%) lags behind humans ($>$95%). In settings with abstract values and country information, model performance drops substantially ($<$60%), while human accuracy remains high ($>$90%). Furthermore, we find that models are better at recognizing socially acceptable versus unacceptable situations. Our findings showcase the current pitfalls in socio-cultural reasoning of LLMs which hinder their adaptability for global audiences.
@@ 1 Introduction
Previous work has largely focused on assessing knowledge and biases by probing LLMs with curated socio-cultural knowledge databases Nguyen et al. (2023); Dwivedi et al. (2023); Fung et al. (2024); Shi et al. (2024), often using direct questions about cultural norms, such as, ‘Is it okay to eat with your left hand in India?”. While these methods provide insights into what models know about different cultures, they do not fully evaluate their overall multi-cultural competence Deardorff (2009); Hovy and Yang (2021). We argue that true cultural competence requires models to not just possess cultural knowledge, but also to apply it flexibly to user-specific scenarios. Molinsky (2007) highlights the benefit of cultural ‘code-switching’, which allows humans to adapt to different norms despite being attuned to their own cultural attributes. Similarly, LLMs should be culturally adaptable, i.e., able to adjust their responses based on the user’s cultural context. By utilizing cultural values provided by or inferred from the user, rather than rigidly adhering to internal biases, these models can ensure effective communication across diverse scenarios.
@@
Figure 1: We introduce NormAd, a framework for testing a language model’s ability to adapt its responses when contextualized with varying levels of cultural information specificity, in contrast to prior methods that directly probe models for their knowledge. We show that LMs struggle to pick up cultural cues when provided with varying levels of context, unlike humans, who can generally recognize such cues.
@@
To address the gap in evaluating the cultural adaptability in LLMs, we introduce the NormAd evaluation framework (§3). Using social norms as a proxy for culture Adilazuarda et al. (2024), NormAd evaluates how models reason about the acceptability of social situations described in free-text, under varying levels of socio-cultural context. As shown in Figure 1, each situational description is evaluated with varying degrees of cultural norm specificity: specific Country names, an abstract high-level Value along with Country names, and fine-grained RoT. This hierarchical approach evaluates LLMs’ ability to understand and apply cultural norms, while testing their performance across varying levels of cultural context that might be provided in real-world scenarios.
@@
As an instantiation of our framework, we develop NormAd-Eti4), a benchmark for measuring cultural adaptability specifically focused on social etiquette norms in English. These multicultural norms are sourced from Cultural Atlas (Evason et al., 2024), an educational resource built on extensive global community interviews and rigorous validation. NormAd-Eti contains 2.6k social situations from 75 countries, each with a question-answer pair to evaluate LLMs’ ability to judge the social acceptability of specific actions across various cultures and levels of cultural norm specificity.
@@ 2.1 Culture in LLMs
Recently, several studies have examined the socio-cultural reasoning of LLMs, evaluating their ability to serve diverse users and values. Some studies have used psychological and cultural surveys WVS (1981); Hofstede (1980) to prompt models with human values Johnson et al. (2022); Atari et al. (2023); Masoud et al. (2023); Ramezani and Xu (2023), guaging how well these models reflect diverse cultural values. Other studies have focused on probing LLMs for their cultural knowledge of social norms (Chiu et al., 2024; Palta and Rudinger, 2023; Shi et al., 2024). While Dwivedi et al. (2023) explored etiquette-related norms through direct probing for knowledge, our approach instead measures adaptability. Studies have also investigated LLMs’ knowledge of cultural artifacts such as food, art forms, clothing, and geographical markers Seth et al. (2024); Li et al. (2024); Koto et al. (2024). These evaluations have primarily focused on measuring the knowledge component of cultures in LLMs, rather than applying and adapting their knowledge to user-specific scenarios. Efforts to improve adaptability have mostly focused on enabling LLMs to adopt synthetic personas from different regions AlKhamissi et al. (2024); Durmus et al. (2023); Kwok et al. (2024).
@@ 2.2 On Value Pluralism and Personalization of LLMs
Cultural studies in LLMs inherently involve dealing with conflicting values, a term known as ‘value pluralism’. Several works have studied this broader problem through either benchmark datasets Ren et al. (2024); Sorensen et al. (2024a); Pistilli et al. (2024), finetuning models to respond pluralistically and prosocially Kim et al. (2022); Forbes et al. (2020) or by proposing modular frameworks around value pluralism Benkler et al. (2023); Feng et al. (2024). Our work is pluralistic in that it prompts LLMs with situations that can have potentially conflicting social acceptabilties depending on context.
@@ 3 NormAd Evaluation Framework
We introduce a multi-level evaluation framework to measure the cultural adaptability of LLMs, contrasting existing work that primarily measures knowledge2.1). To be inclusive of diverse populations with varying values Sorensen et al. (2024b), we argue that a truly adaptable LLM should perform well across diverse cultural contexts.
@@ 4.3 Human Validation
Annotators endorsed our checks’ validity at 84.2% on average, and their interrater agreement yielded a Fleiss fixed-marginal multirater $\kappa=0.56$ and pairwise agreement (PPA) $=0.73$. These results indicate that the annotators overwhelmingly rated our situations, corresponding RoT s, Values, labels, and their relationships as valid, confirming the validity of NormAd. We report per-question scores and payments in Appendix B
(a) RoT
(b) Country +Value
(c) Country
@@ 4.4 Verification of Human Performance
We conduct a small-scale human study considering 3 countries: India, China, South Korea. For each country, we sample 12 data points across 4 subcategories, and select 1 data point per label. We employ 3 graduate students from each country. Averaging across 3 countries, we achieve a Krippendorff’s $\alpha=0.45$ and Fleiss’s $\kappa=0.63$. We compute Value + Country accuracy through majority voting, reporting an overall accuracy of 91.6% against the ground truth labels. The label-wise accuracies are 90% for ‘Yes’, 86.7% for ‘No’, and 91.6% for ‘Neutral’. This highlights that humans from the relevant culture show strong performance at determining the acceptability of situations when conditioned on abstract Value and Country contexts. See App. B.3 for country-wise splits.
@@ 6.1 How well do models perform across different levels of cultural contexts?
LLMs show clear limitations when handling Country and Value +Country contexts, with the best performing models GPT-3.5-turbo, GPT-4footnote 6, and Mistral-7b-Instruct achieving only 59-63% accuracy for Value +Country and 51-56% for Country (see Figure 3). In contrast, our human study across three countries (§4.4) demonstrates that humans can perform very well in these settings, achieving a high accuracy of 90%. The wide performance gap highlights the pressing need for LLMs to better adapt to Country and Value contexts, given that real-world scenarios might often lack specificity wrt cultural cues.
@@
Evaluating the social acceptability under RoT is straightforward since it contains all the necessary information to navigate the specific situation. The QA task essentially reduces to a contextual textual similarity or entailment problem. Our human study (§4.4) demonstrates that humans perform exceptionally well on this task, achieving high 95.6% accuracy. However, models under perform, likely due to a lack of adaptability to textual similarity under cultural and social nuances. The best performing models are GPT-4666 We note that our data was generated with GPT-4, which may give it an unfair advantage; however, even so, we find that GPT-4 still struggles with performance. with 87.6%, Mistral-7b-Instruct with 81.8% and Llama-2-70b-chat with 71.3%, lagging behind human performance. These findings highlight the gap in contextualization capabilities of LLMs, especially with respect to cultural contexts.
@@ 6.2 How well do models perform across the Inglehart-Welzel (IW) cultural map?
We mapped 75 countries into 8 clusters based on the Inglehart-Welzel cultural map. The Country + Value conditioned results, illustrated in Figure 4, show that best-performing models like Llama-2-70b, Llama-1-30b-SFT-KTO, and GPT-4footnote 6 vary in performance across different cultural zones. For instance, they perform better with situations from “English Speaking” countries (e.g., USA) than from “African-Islamic” countries (e.g., Saudi Arabia). In contrast, poorer-performing models, like Llama-2-13b and Llama-1-30b-SFT, under perform consistently across all zones. We hypothesize that larger model sizes and improved training regimes lead to better exploitation of Western cultural cues, causing skewed performance across zones. We see similar trends across Country and RoT (see Appendix D.2). This ‘western-centric’ bias is consistent with model performance on other datasets Johnson et al. (2022); Naous et al. (2023) and observed across various LM architectures Palta and Rudinger (2023) and modalities Ventura et al. (2023).
Figure 4: Model-wise accuracies under Country + Value show a notable performance skew, with top models performing better in social situations from English-speaking countries than in African-Islamic cultural zones.
@@
We find that while DPO and KTO exhibit marginal performance improvements over PPO in the smaller 7b model, their performance significantly improves in the larger 30b model. Figure 5 shows that KTO emerges as the most effective option for cultural adaptability, when conditioned on RoT. We see similar trends for Country and Value + Country as well (see App. D.1).
@@ 6.3 What is the performance across subcategories of NormAd-Eti?
We analyse model performance across the 4 subcategories: ‘Eating’, ‘Gifting’, ‘Visiting’, and ‘Basic Etiquette’. Models consistently underperform in ‘Gifting’, even with RoT conditioning, while they excel in ‘Eating’ and show improved results in ‘Visiting’ and ‘Basic Etiquette’. Our qualitative analysis reveals that ‘Gifting’ involves highly detailed norms regarding the presentation, number, and color of gifts. Further, gift-giving can be highly contextual in some cultures (Stauss, 2023), with differences in expense, presentation, and meaning playing a significant role in societal norms (Hanna and Srivastava, 2015). This highlights the considerable adaptability required from LLMs in handling such complex social norms. Table 6 in App. D.4 presents some failure cases of GPT-3.5-turbo.
@@ 6.4 How well do models do across social acceptabilities (Yes/No/Neutral)?
We analyze how the social acceptability labels of situations affects model performance. Figure 6 shows the averaged label-wise accuracies of our overall best-performing models (Llama-2-70b, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 6, Llama-1-30b-SFT-KTO). Models generally perform better on situations that conform to social norms (labeled ‘Yes’), and improve on norm-violating situations (labeled ‘No’) with increasing levels of context, indicating that inherent agreement biases within LLMs could impact their adaptability (Sun et al., 2024; Perez et al., 2022).
@@
Figure 6: Averaged accuracy of best performing models (Llama-2-70b, Llama-1-30b-SFT-KTO, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 6) across ground-truth labels. Models are biased towards “yes” (i.e conformations) and worse at “no” (i.e. violations) and “neutral” (i.e irrelevant).
@@ 8 Limitations
Our research examines LLM’s ability to adapt to cultural nuances through a test bed of social situations. However, certain limitations present in our dataset and evaluation framework may warrant further research, such as:
@@
Adilazuarda et al. (2024) acknowledges the difficulty of defining what ‘culture’ means in the context of language models, and categorize prior work according to cultural proxies, linguistic interactions and measurement strategies. NormAd can be thus categorized as a black-box LM evaluation employing a semantic proxy to test models’ reasoning abilities concerning etiquette-related social norms (§4.1), with analyses on demographically informed axes (§6.2). While our approach provides valuable insights on LLM’s cultural competencies, we believe that a more comprehensive picture of LLMs’ cultural landscape can be painted through other demographically and lingusitcally informed proxies and frameworks, and , by broadening the scope of evaluation Bhatt and Diaz (2024).
@@
We conduct evaluations only in English, as Shen et al. (2024) highlight that prompting in English helps models leverage knowledge more effectively and mitigates issues arising from their varying linguistic capabilities and instruction-following skills. However, we acknowledge the significance of multilingualism and linguistic variations such as AAVE (African American Vernacular English) in ethical and cultural studies of AI (Li et al., 2024; Sap et al., 2019; Naous et al., 2023). Our study excludes these aspects from the evaluation due to the extensive validation efforts required for curating a synthetically generated and translated test bed across all languages and linguistic variations. This points to an area for future investigation for cultural adaptability studies, correlating multilingual reasoning performance with reasoning over cultures associated with such languages.
Granularity of Cultural Norms:
While our curated test bed encompasses narratives from 75 countries, it is essential to note the existence of more than 75 countries globally, each characterized by unique and finely-grained cultural norms. Additionally, there can be multiple cultural variations within a single country. We opted to scope our study based on the Cultural Atlas as a verified source of information, recognizing that certain nuances may not be fully captured within this framework.
@@
Most current work in ‘culture’, including our own through NormAd-Eti, often treats the dynamic and multifaceted nature of culture as static variables during evaluations. This static approach may inadvertently perpetuate cultural stereotypes and fail to capture the continuously changing cultural nuances. To address these limitations, we suggest a modification to our evaluation framework, envisioned as future work, that would allow users to specify their own norms and values. Our framework, NormAd, is designed to be flexible, enabling these modifications easily, which is crucial for accommodating the evolving nature of cultural contexts. Improving our framework in this manner would greatly benefit the adaptability and cultural sensitivity of LLMs, ensuring they remain relevant and effective across diverse global environments.
@@
In this work, we study the cultural adaptability of LLMs – can LLMs align to human values across varying cultural contexts? While we suggest working on improving LLMs capabilities on this front, we acknowledge the existence of several human-computer interaction studies that assess the impact of personifying language models to cater to multiple demographics such as Black Americans (Harrington and Egede, 2023). These studies show that identity-related concepts such as age and racial likeness play a part in maintaining user trust and comfort. However, care must be taken when constructing personalized LLMs, owing to societal-risks such as the propagation of polarized views between historically conflicting demographics, and disparities in access among these subpopulations (Kirk et al., 2023). Finally, while we maintain equality over representing all demographics and cultures, we do not discuss the impact of the interaction and access disparities between these subpopulations on the use of language technologies, and the lack of representation they may cause.

### v8

@@ Abstract
To be effectively and safely deployed to global user populations, large language models (LLMs) may need to adapt outputs to user values and cultures, not just know about them. We introduce NormAd, an evaluation framework to assess LLMs’ cultural adaptability, specifically measuring their ability to judge social acceptability across varying levels of cultural norm specificity, from abstract values to explicit social norms. As an instantiation of our framework, we create NormAd-Eti, a benchmark of 2.6k situational descriptions representing social-etiquette related cultural norms from 75 countries. Through comprehensive experiments on NormAd-Eti, we find that LLMs struggle to accurately judge social acceptability across these varying degrees of cultural contexts and show stronger adaptability to English-centric cultures over those from the Global South. Even in the simplest setting where the relevant social norms are provided, the best LLMs’ performance (<82%) lags behind humans (>95%). In settings with abstract values and country information, model performance drops substantially (<60%), while human accuracy remains high (>90%). Furthermore, we find that models are better at recognizing socially acceptable versus unacceptable situations. Our findings showcase the current pitfalls in socio-cultural reasoning of LLMs which hinder their adaptability for global audiences.111 We release the dataset on GitHub here: https://github.com/Akhila-Yerukola/NormAd
@@ 1 Introduction
Previous work has largely focused on assessing knowledge and biases by probing LLMs with curated socio-cultural knowledge databases Nguyen et al. (2023); Dwivedi et al. (2023); Fung et al. (2024); Shi et al. (2024), often using direct questions about cultural norms, such as, “Is it okay to eat with your left hand in India?”. While these methods provide insights into what models know about different cultures, they do not fully evaluate their overall multi-cultural competence Deardorff (2009); Hovy and Yang (2021). We argue that true cultural competence requires models to not just possess cultural knowledge, but also to apply it flexibly to user-specific scenarios. Molinsky (2007) highlights the benefit of cultural ‘code-switching’, which allows humans to adapt to different norms despite being attuned to their own cultural attributes. Similarly, LLMs should be culturally adaptable Chang et al. (2013), i.e., able to adjust their responses based on the user’s cultural context. While it is still an open question as to how quickly or to what extent LLMs need to be adaptable, they can ensure effective communication across diverse scenarios by utilizing cultural values provided by or inferred from the user, rather than rigidly adhering to internal biases.
@@
Figure 1: We introduce NormAd, a framework for testing a language model’s ability to adapt its responses when contextualized with varying levels of cultural information specificity, in contrast to prior methods that directly probe models for their knowledge. We show that LMs struggle to pick up cultural cues when provided with varying levels of context (Xs representing their incorrect responses, unlike humans, who can generally recognize such cues.)
@@
To address the gap in evaluating the cultural adaptability in LLMs, we introduce the NormAd evaluation framework (§3). Using social norms as a proxy for culture Adilazuarda et al. (2024), NormAd evaluates how models reason about the acceptability of social situations described in free-text, under varying levels of socio-cultural context. As shown in Figure 1, each situational description is evaluated with varying degrees of cultural norm specificity: specific Country names, an abstract high-level Value along with Country names, and fine-grained Rules-of-Thumb. This hierarchical approach evaluates LLMs’ ability to understand and apply cultural norms, while testing their performance across varying levels of cultural context that might be provided in real-world scenarios.
@@
As an instantiation of our framework, we develop NormAd-Eti4), a benchmark for measuring cultural adaptability specifically focused on social etiquette norms specified in English. These multicultural norms are sourced from the Cultural Atlas (Evason et al., 2024), an educational resource based on extensive global community interviews and rigorous validation. NormAd-Eti contains 2.6k descriptions of social situations from 75 countries, each with a question-answer pair to evaluate LLMs’ ability to judge the social acceptability of specific actions across various cultures and levels of cultural norm specificity.
@@ 2.1 Culture in LLMs
Recently, several studies have examined the socio-cultural reasoning of LLMs, evaluating their ability to serve diverse users and values. Some studies have used psychological and cultural surveys WVS (1981); Hofstede (1980) to prompt models with human values Johnson et al. (2022); Atari et al. (2023); Masoud et al. (2023); Ramezani and Xu (2023), gauging how well these models reflect diverse cultural values. Other studies have focused on probing LLMs for their cultural knowledge of social norms (Chiu et al., 2024; Palta and Rudinger, 2023; Shi et al., 2024). While Dwivedi et al. (2023) explored etiquette-related norms through direct probing for knowledge, our approach instead measures adaptability. Studies have also investigated LLMs’ knowledge of cultural artifacts such as food, art forms, clothing, and geographical markers Seth et al. (2024); Li et al. (2024); Koto et al. (2024). These evaluations have primarily focused on measuring the knowledge component of cultures in LLMs, rather than applying and adapting their knowledge to user-specific scenarios. Efforts to improve adaptability have mostly focused on enabling LLMs to adopt synthetic personas from different regions AlKhamissi et al. (2024); Durmus et al. (2023); Kwok et al. (2024).
@@ 2.2 On Value Pluralism and Personalization of LLMs
Cultural studies in LLMs inherently involve dealing with conflicting values, a term known as ‘value pluralism’. Several works have studied this broader problem through either benchmark datasets Ren et al. (2024); Sorensen et al. (2024a); Pistilli et al. (2024), finetuning models to respond pluralistically and prosocially Kim et al. (2022); Forbes et al. (2020) or by proposing modular frameworks around value pluralism Benkler et al. (2023); Feng et al. (2024). Our work is pluralistic in that it prompts LLMs with situations that can have potentially conflicting social acceptabilities depending on context.
@@ 3 NormAd Evaluation Framework
We introduce a multi-level evaluation framework to measure the cultural adaptability of LLMs, contrasting existing work that primarily measures knowledge2.1). Borrowing from Chang et al. (2013), we say that an LLM is culturally adaptable if its outputs are personalized or adapted towards multicultural users. [^3] We make a distinction from cultural adaptation/transcreation Nida (1964), which involves adapting an aspect of one culture to another. To be inclusive of diverse populations with varying values Sorensen et al. (2024b), we argue that a truly adaptable LLM should perform well across diverse user-provided cultural contexts Varshney (2023).
@@ 4.3 Human Validation
Annotators endorsed our checks’ validity at 84.2% on average, and their interrater agreement yielded a Fleiss fixed-marginal multirater $\kappa=0.56$ and pairwise agreement (PPA) $=0.73$. These results indicate that the annotators overwhelmingly rated our situations, corresponding RoT s, Values, labels, and their relationships as valid, confirming the validity of NormAd. We report per-question scores and payments in Appendix B.
@@ 4.4 Verification of Human Performance
We conduct a small-scale human study considering 3 countries: India, China, South Korea. For each country, we sample 12 data points across 4 subcategories, and select 1 data point per label. We employ 3 graduate students from each country. Averaging across 3 countries, we achieve a Krippendorff’s $\alpha=0.45$ and Fleiss’s $\kappa=0.63$. We compute Value + Country accuracy through majority voting, reporting an overall accuracy of 91.6% against the ground truth labels. The label-wise accuracies are 90% for ‘Yes’, 86.7% for ‘No’, and 91.6% for ‘Neutral’. This highlights that humans from the relevant culture show strong performance at determining the acceptability of situations when conditioned on abstract Value and Country contexts. Please refer to Appendix B.3 for country-wise splits.
(a) RoT
(b) Country +Value
(c) Country
@@ 5 Experimental Setup
We evaluate several language models on their ability to adapt to varying levels of cultural contexts.
@@ 6.1 How well do models perform across different levels of cultural contexts?
LLMs show clear limitations when handling Country and Value +Country contexts, with the best performing models GPT-3.5-turbo, GPT-4footnote 8, and Mistral-7b-Instruct achieving only 59-63% accuracy for Value +Country and 51-56% for Country (see Figure 3). In contrast, our human study across three countries (§4.4) demonstrates that humans can perform very well in these settings, achieving a high accuracy of 90%. The wide performance gap highlights the pressing need for LLMs to better adapt to Country and Value contexts, given that real-world scenarios might often lack specificity wrt cultural cues.
@@
Evaluating the social acceptability under RoT is straightforward since it contains all the necessary information to navigate the specific situation. The QA task essentially reduces to a contextual textual similarity or entailment problem. Our human study (§4.4) demonstrates that humans perform exceptionally well on this task, achieving high 95.6% accuracy. However, models under perform, as shown in Figure 3, likely due to a lack of adaptability to cultural and social nuances in textual similarity tasks. The best performing models are GPT-4888 We note that our data was generated with GPT-4, which may give it an unfair advantage; however, even so, we find that GPT-4 still struggles with performance. with 87.6%, Mistral-7b-Instruct with 81.8% and Llama-2-70b-chat with 71.3%, lagging behind human performance. These findings highlight the gap in contextualization capabilities of LLMs, especially with respect to cultural contexts.
@@
Figure 4: Comparision of model accuracies under Country + Value shows a notable performance skew, with top models (with increased size or improved preference alignment methods) performing better in social situations from English-speaking countries than in African-Islamic cultural regions.
@@ 6.2 How well do models perform across the Inglehart-Welzel (IW) cultural map?
We mapped 75 countries into 8 clusters based on the Inglehart-Welzel cultural map. The Country + Value conditioned results, illustrated in Figure 4, show that best-performing models like Llama-2-70b, Llama-1-30b-SFT-KTO, and GPT-4footnote 8 vary in performance across different cultural zones. For instance, they perform better with situations from “English Speaking” countries (e.g., USA) than from “African-Islamic” countries (e.g., Saudi Arabia). In contrast, poorer-performing models, like Llama-2-13b and Llama-1-30b-SFT, under perform consistently across all zones. We hypothesize that larger model sizes and improved training regimes lead to better exploitation of Western cultural cues, causing skewed performance across zones. We see similar trends across Country and RoT (see Appendix D.2). This ‘western-centric’ bias is consistent with model performance on other datasets Johnson et al. (2022); Naous et al. (2023) and observed across various LM architectures Palta and Rudinger (2023) and modalities Ventura et al. (2023).
@@
We find that while DPO and KTO exhibit marginal performance improvements over PPO in the smaller 7b model, their performance significantly improves in the larger 30b model. Figure 5 shows that KTO emerges as the most effective option for cultural adaptability, when conditioned on RoT. We see similar trends for Country and Value + Country as well (see Appendix D.1 for more details).
Figure 6: Averaged accuracy of best performing models (Llama-2-70b, Llama-1-30b-SFT-KTO, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 8) across ground-truth labels. Models are biased towards “yes” (i.e conformations) and worse at “no” (i.e. violations) and “neutral” (i.e irrelevant).
@@ 6.3 What is the performance across subcategories of NormAd-Eti?
We analyse model performance across the 4 subcategories: ‘Eating’, ‘Gifting’, ‘Visiting’, and ‘Basic Etiquette’. Models consistently underperform in ‘Gifting’, even with RoT conditioning, while they excel in ‘Eating’ and show improved results in ‘Visiting’ and ‘Basic Etiquette’. Our qualitative analysis reveals that ‘Gifting’ involves highly detailed norms regarding the presentation, number, and color of gifts. Further, gift-giving can be highly contextual in some cultures (Stauss, 2023), with differences in expense, presentation, and meaning playing a significant role in societal norms (Hanna and Srivastava, 2015). We additionally present our quantitative findings for subcategories in Figure 11 in Appendix D. Most models exhibit a performance dip for the ‘gifts’ subaxis. The Country + Value/ RoT contextualization mitigates this drop to some extent for some (but not all) models. This highlights the considerable adaptability required from LLMs in handling such complex social norms. Table 7 in Appendix D.4 presents some failure cases of GPT-3.5-turbo.
@@ 6.4 How well do models do across social acceptabilities (Yes/No/Neutral)?
We analyze how the social acceptability labels of situations affects model performance. Figure 6 shows the averaged label-wise accuracies of our overall best-performing models (Llama-2-70b, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 8, Llama-1-30b-SFT-KTO). Models generally perform better on situations that conform to social norms (labeled ‘Yes’), and improve on norm-violating situations (labeled ‘No’) with increasing levels of context, indicating that inherent agreement biases within LLMs could impact their adaptability (Sun et al., 2024; Perez et al., 2022).
@@ 8 Limitations
Our research examines LLMs’ abilities to adapt to cultural nuances through a test bed of social situations. However, certain limitations present in our dataset and evaluation framework may warrant further research, such as:
@@
Defining ‘culture’, especially in the context of language models is challenging, with prior work categorizing approaches by cultural proxies, linguistic interactions, and measurement strategies Adilazuarda et al. (2024). NormAd employs a black-box evaluation approach, using etiquette-related social norms as a semantic proxy of culture, with analyses on demographically informed axes (§6.2). While this approach offers valuable insights into LLMs’ cultural competencies, broader evaluation through diverse proxies is needed to capture the full cultural landscape (Bhatt and Diaz, 2024).
Cultural Diversity and Representation:
Cultural norms are highly diverse, with significant variation within countries, across regions, and among social classes. The Cultural Atlas only captures the dominant cultural narrative present in each country, leaving several variations unrepresented. Future work should should build resources that capture these diverse cultural perspectives and evaluate models on their ability to adapt across them.
@@
In this work, we conduct evaluations only in English. While testing across multiple languages and linguistic variations is essential for robust benchmarking of LLMs, it is beyond the scope of this study. Prior work highlights that prompting in English – given current skewed data representations – helps models leverage knowledge more effectively and mitigates issues arising from varying linguistic capabilities and instruction-following skills Shen et al. (2024). We encourage future work to investigate multilingual reasoning performance and its correlation with cultural adaptability across languages.
@@
As a pragmatic way for approaching culture, much research, including our own through NormAd-Eti, often treats the dynamic and multifaceted nature of culture as static variables during evaluation. This static approach may inadvertently perpetuate cultural stereotypes and fail to capture the continuously changing cultural nuances. To address these limitations, we suggest a modification to our evaluation framework, envisioned as future work, that would allow users to specify their own norms and values. Our framework, NormAd, is designed to be flexible, which is crucial for accommodating the evolving nature of cultural contexts.
@@
In this work, we study the cultural adaptability of LLMs – specifically, can LLMs align to human values across varying cultural contexts? While we advocate for improving LLM capabilities in this area, we recognize the complexities involved. Prior human-computer interaction studies suggest that personifying language models to cater to multiple demographics, such as Black Americans, can enhance user trust and comfort (Harrington and Egede, 2023; Wenzel and Kaufman, 2024). However, the extent to which LLMs should adapt to users’ cultural nuances remains uncertain. Excessive adaptation risks mimicry that may be perceived as manipulative, undermining user trust, particularly if the adaptation is seen as a shortcut to gaining social acceptance within a subgroup. Moreover, highly adaptable systems may amplify societal risks, such as reinforcing polarized views between historically conflicting demographics (Kirk et al., 2023). These complexities are further compounded by the fact that cultural norms are not monolithic; multiple variations often exist within a single country, region, or social group. Addressing this diversity requires adaptable frameworks that empower users to prescribe their cultural values, or opt out of certain adaptations altogether. Crucially, LLMs should adapt based on user-provided preferences rather than impose cultural norms. As a first step towards this, NormAd provides a framework for measuring cultural adaptability, with our benchmark NormAd-Eti merely serving as a proxy for measuring adaptability rather than prescribing cultural standards.
Acknowledgements
This project is funded in part by NSF grant (#2230466), DSO National Laboratories, and an Amazon Research Award, Spring 2023 CFP. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of Amazon. We additionally acknowledge the Cultural Atlas, a social cohesion project delivered by Mosaica and SBS, and the team that curated its cultural values, norms, and specifications. Finally, we would like to thank Shaily Bhatt and Joel Mire for their insightful comments on our work.